# Benchmarking pig detection and tracking under diverse and challenging conditions

### v1

@@ Acknowledgements
During the preparation of this work, the authors used ChatGPT [102] in order to paraphrase text and generate basic code for data processing as well as figure and table generation. After using ChatGPT, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A Detection datasets
Alameer [51] 630×353 811×509 308 153 3278 75 2a
Bergamini [10] 1280×720 1280×720 186 0 1801 36 1
Psota [50] 1920×1080 3840×2160 733 91 10 3011030110\,30110 301 141 17
Henrich (our) 640×360 1920×1080 1704 443 26 7972679726\,79726 797 316 11
@@
Figure 7: Example images from all unique pen environments contained in the publicly available annotated object detection datasets presented in Sec. 2.2. The character in the top left corner of each image indicates the initial of the first author of the corresponding work (A[5], M[30], R[11, 25]). The unique pen environments presented in each work are then numbered starting from one.
Figure 8: Example images from all unique pen environments contained in the PigDetect dataset. The character in the top left corner of each image indicates the initial of the first author of the work in which the original image data was released (B[10], A[51], P[50]). H denotes Henrich, representing data introduced in this work. The unique pen environments from each source are then numbered starting from one. H6 is the pen environment of the PigDetect test set. Apart from different pen environments, PigDetect includes images from different camera angles. These are not depicted here.
@@ Appendix B Tracking datasets
See Tab. 10 as well as Fig. 9 and Fig. 10 on the following two pages.
Figure 9: Example images from all unique pen environments contained in the publicly available annotated multi-object tracking datasets presented in Sec. 3.2. The character in the top left corner of each image indicates the initial of the first author of the corresponding work (B[10], M[30], Y[42], J[40]). The unique pen environments presented in each work are then numbered starting from one.

### v2

@@ Acknowledgements
During the preparation of this work, the authors used ChatGPT [OpenAI2023] in order to paraphrase text and generate basic code for data processing as well as figure and table generation. After using ChatGPT, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A Detection datasets
Alameer [alameer_automated_2022] 630×353 811×509 308 153 3278 75 2a
Bergamini [bergamini_extracting_2021] 1280×720 1280×720 186 0 1801 36 1
Psota [psota_multi-pig_2019] 1920×1080 3840×2160 733 91 10 30110\,301 141 17
Henrich (our) 640×360 1920×1080 1704 443 26 79726\,797 316 11
@@
Figure 7: Example images from all unique pen environments contained in the publicly available annotated object detection datasets presented in Sec. 2.2. The character in the top left corner of each image indicates the initial of the first author of the corresponding work (A[alameer_automated_2020], M[melfsen2023describing], R[riekert_automatically_2020, riekert2021model]). The unique pen environments presented in each work are then numbered starting from one.
Figure 8: Example images from all unique pen environments contained in the PigDetect dataset. The character in the top left corner of each image indicates the initial of the first author of the work in which the original image data was released (B[bergamini_extracting_2021], A[alameer_automated_2022], P[psota_multi-pig_2019]). H denotes Henrich, representing data introduced in this work. The unique pen environments from each source are then numbered starting from one. H6 is the pen environment of the PigDetect test set. Apart from different pen environments, PigDetect includes images from different camera angles. These are not depicted here.
Figure 9: Examples for near-identical training and test images in the detection dataset published by melfsen2023describing. For reference, we also provide the names of the respective images in the published dataset: Train 1 (954-0000-0001_400.jpg), Train 2 (836-0125-0126_750.jpg), Train 3 (931-0035-0037_3150.jpg), Train 4 (ch01_20211120121240.mp4_10.jpg), Test 1 (954-0000-0001_450.jpg), Test 2 (836-0125-0126_800.jpg), Test 3 (931-0035-0037_3100.jpg), Test 4 (ch01_20211120121240.mp4_9.jpg).
Figure 10: Visualization of ground truth and predicted bounding boxes for three pens that were not part of the training dataset. Top shows an example from the PigDetect test set, middle a frame from the test video published by yu2025fto, and bottom an image from the object detection test set published by melfsen2023describing. All predicted bounding boxes of the Co-DINO object detector (see Sec. 4.1.3) with a score $\geq 0.2$ are visualized. The three images were selected for visualization because they depict challenging scenarios with poor visibility, piling and occlusions. The detection results were not examined prior to selecting the images. For reference, we also provide the names of the respective images in the published datasets: PigDetect (danuma_1596.jpg), melfsen2023describing (931-0035-0037_3100.jpg), yu2025fto (jochiwon5M_00000865.jpg). The video frame from yu2025fto was rotated $90^{\circ}$ clockwise to better fit the page layout.
@@ Appendix B Tracking datasets
See Tab. 11 and Fig. 11 and 12 on the next two pages.
Figure 11: Example images from all unique pen environments contained in the publicly available annotated multi-object tracking datasets presented in Sec. 3.2. The character in the top left corner of each image indicates the initial of the first author of the corresponding work (B[bergamini_extracting_2021], M[melfsen2023describing], Y[yu2025fto], J[jaoukaew2024robust]). The unique pen environments presented in each work are then numbered starting from one.